OpenStudy (anonymous):

The equation of line AB is x=1/2y +1 . Line CD contains points C(6,5) and D(7,3) . Are the Lines AB and CD parallel, perpendicular or oblique?

OpenStudy (anonymous):

Find the slope of line CD. To do this, you do (y2-y1)/(x2-x1). For line CD, this would be (3-5)/(7-6) which equals (-2)/1 which equals -2. For lines to be parallel, their slopes are opposite reciprocals, meaning the sign is changed and the slope is divided into 1. Slope AB is 1/2, and Slope CD is -2, so this fits, and the lines are perpendicular.

OpenStudy (anonymous):

@sophflake the slope of AB is not 1/2, reread the expression, it says x = 1/2 y +1, not y = 1/2 x +1 !!!

OpenStudy (anonymous):

Oh whoops. Well then, x=1/2y+1 becomes 1/2/y=x-1, becomes y=2x-2. Therefore, the lines are neither parallel or perpendicular, and are therefore oblique.